# Bias in Decision-Making for AI’s Ethical Dilemmas: A Comparative Study of Open-source and Closed-source Large Language Models

### v3

@@ Introduction
For the single protected attribute scenario, each attribute was mentioned in the question. The normalized frequency is the frequency at which a protected attribute is selected under the requisitions for which the LLMs choose a protected attribute in the protected attribute group. For example, in $G_{Gender}$ all protected attributes is selected 8 times, and the protected attribute $Masculine$ is selected 4 times, and the normalized frequency of $Masculine$ is 0.5.
@@
where $f_{pa}$ is the normalized frequency for protected attribute $pa$ of category $G$, $N_{pa}$ is the count of $pa$ appeared in the experiment. For example, $f_{Masculine}$ is the normalized frequency for protected attribute $Masculine$ of category $G_{Gender}$, $N_{Masculine}$ is the count of $Masculine$ appeared in the experiment and is 5. The $\sum_{pa\in G_{Gender}}N_{pa}$ is 10. And the normalized frequency for protected attribute $Masculine$ of category $G_{Gender}$ is 0.5.
@@
where $f^{*}_{k}$ is the mean normalized frequency of the specific single protected attribute $k$ in intersectional scenario, $Count_{pa_{1}=k}$ is the number of the intersectional protected attributes including the specific single protected attribute $k$. For example, $f^{*}_{Masculine}$ is the mean normalized frequency of the specific single protected attribute $Masculine$ in intersectional scenario, $Count_{pa_{1}=Masculine}$ is the number of the intersectional protected attributes including the specific single protected attribute $Masculine$ and is 17, the $\sum_{pa_{1}=k}f_{pa_{1},pa_{2}}$ is 3.4, and the mean normalized frequency of the specific single protected attribute $Masculine$ in intersectional scenario is 0.2.
@@
For single protected attribute scenario, we directly used the standard deviation of the normalized frequency of each protected attribute to assess the stability of the ethical preferences of LLMs. We designed the ethical stability as the normalized total standard deviation of the protected attribute group. For example, the normalized total standard deviation of the specific single protected attribute group $Gender$ is 0.2, and the ethical stability of the protected attribute group $Gender$ is 0.2. The smaller the standard deviation, the more stable the ethical preference.
@@
where $\mu_{a}$ is the cluster’s mean value, and $f^{*}_{pa}$ is the mean normalized frequency of protected attribute $pa$, $n_{a}$ is the number of samples in cluster $a$.
@@
where $B_{pa}\in[-1,1]$ is the preference score of protected attribute $pa$, $f_{pa}^{Open}$ and $f_{pa}^{Closed}$ are the mean normalized frequencies of protected attribute $pa$ for Open-source LLMs and Closed-source LLMs. For example, $B_{Black}$ is the preference score of protected attribute $Black$, $f_{Black}^{Open}$ and $f_{Black}^{Closed}$ are the mean normalized frequencies of protected attribute $Black$ for Open-source LLMs and Closed-source LLMs and are 0.5 and 0.3, and the preference score $B_{Black}$ is 0.25.

### v5

@@ Introduction
For the single protected attribute scenario, each attribute was mentioned in the question. The normalized frequency is the frequency at which a protected attribute is selected under the requisitions for which the LLMs choose a protected attribute in the protected attribute group. For example, in $G_{Gender}$ all protected attributes is selected 8 times, and the protected attribute MasculineMasculine is selected 4 times, and the normalized frequency of MasculineMasculine is 0.5.
@@
where $f_{pa}$ is the normalized frequency for protected attribute papa of category $G$, $N_{pa}$ is the count of papa appeared in the experiment. For example, $f_{Masculine}$ is the normalized frequency for protected attribute MasculineMasculine of category $G_{Gender}$, $N_{Masculine}$ is the count of MasculineMasculine appeared in the experiment and is 5. The $\sum_{pa\in G_{Gender}}N_{pa}$ is 10. And the normalized frequency for protected attribute MasculineMasculine of category $G_{Gender}$ is 0.5.
@@
where $f^{*}_{k}$ is the mean normalized frequency of the specific single protected attribute $k$ in intersectional scenario, $Count_{pa_{1}=k}$ is the number of the intersectional protected attributes including the specific single protected attribute $k$. For example, $f^{*}_{Masculine}$ is the mean normalized frequency of the specific single protected attribute MasculineMasculine in intersectional scenario, $Count_{pa_{1}=Masculine}$ is the number of the intersectional protected attributes including the specific single protected attribute MasculineMasculine and is 17, the $\sum_{pa_{1}=k}f_{pa_{1},pa_{2}}$ is 3.4, and the mean normalized frequency of the specific single protected attribute MasculineMasculine in intersectional scenario is 0.2.
@@
For single protected attribute scenario, we directly used the standard deviation of the normalized frequency of each protected attribute to assess the stability of the ethical preferences of LLMs. We designed the ethical stability as the normalized total standard deviation of the protected attribute group. For example, the normalized total standard deviation of the specific single protected attribute group GenderGender is 0.2, and the ethical stability of the protected attribute group GenderGender is 0.2. The smaller the standard deviation, the more stable the ethical preference.
@@
where $\mu_{a}$ is the cluster’s mean value, and $f^{*}_{pa}$ is the mean normalized frequency of protected attribute papa, $n_{a}$ is the number of samples in cluster $a$.
@@
where $B_{pa}\in[-1,1]$ is the preference score of protected attribute papa, $f_{pa}^{Open}$ and $f_{pa}^{Closed}$ are the mean normalized frequencies of protected attribute papa for Open-source LLMs and Closed-source LLMs. For example, $B_{Black}$ is the preference score of protected attribute BlackBlack, $f_{Black}^{Open}$ and $f_{Black}^{Closed}$ are the mean normalized frequencies of protected attribute BlackBlack for Open-source LLMs and Closed-source LLMs and are 0.5 and 0.3, and the preference score $B_{Black}$ is 0.25.